OpenStudy (anonymous):

someone please help.... Find the value of x: 2^1/x +1= 3.3846

OpenStudy (anonymous):

2^(1/x)+1=3.3846 2^(1/x)=2.3846 ln(2^(1/x))=ln(2.3846) ln(2)/x=ln(2.3846) x=ln(2)/ln(2.3846)

OpenStudy (anonymous):

\[\approx .798\]

OpenStudy (anonymous):

i dont understand how u got to this part: ln(2)/x=ln(2.3846) then..x=ln(2)/ln(2.3846)

OpenStudy (anonymous):

2^(1/x)+1=3.3846 2^(1/x)=2.3846 ln(2^(1/x))=ln(2.3846) multiple both side by ln ln(2)/x=ln(2.3846) property of logarithms, the exponent goes to the front -> 1/x ln(2) rewrite -> ln(2)/x x=ln(2)/ln(2.3846) solving for x

OpenStudy (anonymous):

thanks